Daniel spent $9 on 1 daylily and 1 geranium. Scott spent $42 on 4 daylilies and 5 geraniums. What is the cost of 1 daylily and the cost of one geranium?
Answer by richwmiller(17219)

1d+1g=9
4d+5g=42
4d+4g=36
g=6
d=3
check
4*3=12
5*6=30
30+12=42